# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v3

@@ III Revisiting QRNG based on Single-photon Source
Figure 1: A generic block diagram of the single-photon-based QRNG. Single-photon, coming from the source, are sent through a $/50$ 50 beam-splitter. Two SPDs ($D_{0},D_{1}$) are placed in two paths to detect the photon. Photon may take either path with probability $0.5$. The random bit would be $i$ if the detection happens at $D_{i}$.
If the photon source for this QRNG is not a perfect single-photon source, it may happen that at the same time some photons are transmitted through the beam-splitter and some are reflected. Then, both the detectors may detect photons simultaneously, and the random number cannot be generated. To ensure whether the source is a single-photon source or not, one may perform the first step of the two-fold statistical method mentioned in Section V. If this method reveals that the distribution is not sub-Poissonian, then the source cannot be a single-photon source, and that source cannot be used in this QRNG.
Each bit of the random number generated with this single-photon-based QRNG has probability $\frac{1}{2}$ of being 00 or $1$. So, ideally it produces uniform random numbers. However, any defect in the beam-splitter or the detectors may change this probability. For example, consider an $R/T$ beam-splitter and the detectors $D_{i}$ with detection efficiency $d_{i}\in(0,1]$. Suppose, $D_{0}$ is placed on the path of the transmitted light, and $D_{1}$ is placed on the path of the reflected light. Then the probability of 00 in the generated random number, that is, the probability of detecting a photon by detector $D_{0}$ would be
@@
If $d_{0}T\neq d_{1}R$, the generated random number would not be uniformly random. To test this randomness, statistical tests like NIST, Dieharder, AIS-$31$, ENT etc. may be used. See Section VIII for a discussion on these.
IV Revisiting QRNG based on Coherent Pulse
It is well known that producing single photon states is experimentally challenging Imamoḡlu and Yamamoto (1994); Kim et al. (1999). However, different experimental approaches Brunel et al. (1999); Lounis and Moerner (2000); Santori et al. (2001, 2002); Vučković et al. (2003); Kuhn et al. (2002); Adam and Mechler (2024); Wang and Wang (2024) have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval Argillander et al. (2023). In this process, the photons follow Poissonian statistics. A generic block diagram of the photon arrival time-based QRNG using photons from coherent sources is given in Fig. 2.
@@ IV Revisiting QRNG based on Coherent Pulse
Figure 2: A basic block diagram of the photon arrival time-based QRNG using weak coherent pulses. Coherent pulses, coming from the source, are sent through attenuators to reduce their power so that the probability of having multiple photons within a predefined time interval becomes negligible Argillander et al. (2023). A SPD detects these photons and the detection time (photon arrival time) has been registered. The Field Programmable Gate Array (FPGA) provides random numbers based on the registered time stamps.
@@
There are different methods to produce random numbers from the registered times. Time diagrams for these methods are shown in Fig. 3. In Fig. 3(I), if the registered times of photon arrivals are $\{t_{0},t_{1},\dots,t_{2i-2},t_{2i-1},t_{2i},\dots\}$, the non-overlapping consecutive time intervals, i.e., $\Delta t_{2i-1}=t_{2i-1}-t_{2i-2},\Delta t_{2i}=t_{2i}-t_{2i-1}$, are used to generate random numbers as follows.
$\begin{cases}\text{The $i$-th bit is taken as $0$, if }\Delta t_{2i-1}<\Delta t_{2i},\\ \text{The $i$-th bit is taken as $1$, if }\Delta t_{2i-1}>\Delta t_{2i}.\end{cases}$ (9)
But in this process Dynes et al. (2008), the random number generation rate is very slow. The length of the random bit-string is half of the photon count. Comparatively faster processes have been developed in the literature Stipčević and Rogina (2007); Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010); Wahl et al. (2011); Nie et al. (2014, 2014). In Fig. 3(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i-1$ is taken as the random number. A new cycle of the bins starts after the detector has a fixed sleep time. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by
@@
Therefore, the random number generated with this procedure follows exponential distribution Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010); Wahl et al. (2011); Nie et al. (2014). Another QRNG scheme is proposed in Stipčević and Rogina (2007), as depicted in Fig. 3(III), using external time reference. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i-1$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
@@
V Our Proposal of Two-fold Statistical Method for Testing Quantumness
In Section II.1 we have discussed that for the Poissonian distribution, the mean of the photon-counting distribution is the same as its variance. However, it is interesting to note that mean $=$ variance does not always imply that the distribution is Poisson. For example, consider the probability distribution
@@ V Our Proposal of Two-fold Statistical Method for Testing Quantumness
This motivates us to propose a two-fold method to test the quantumness of the photon source as an answer to question Q1. First, a statistical test is performed to verify whether the mean is equal to or less than or greater than the variance. Note that any statistical validation of equality is not exact, but rather approximate in practice. Thus, if the above test validates the mean and the variance to be equal, we propose to perform a second test to validate that the distribution is indeed Poisson.
V.1 Interval Estimation
Since Poissonian and super-Poissonian statistics are explainable using classical theory, only sub-Poissonian distribution can guarantee the quantumness of a quantum device. We suggest a procedure to decide whether the underlying distribution is sup-Poissonian or not by estimating the confidence interval for the mean of the underlying distribution as follows.
Suppose, the number of photon counts within a fixed time is given as a sample of size $n$. Then by the central limit theorem, $\frac{\langle X\rangle-m}{\sigma/\sqrt{n}}$ is approximately standard normal for large $n$  Hogg and Craig (2019), where $\langle X\rangle$ is the random variable corresponding to the sample mean $\langle x\rangle$, and $m,\sigma^{2}$ are mean and variance of the actual photon-count distribution. Let the sample variance be given by $S^{2}$. Then $s^{2}=\frac{n}{n-1}S^{2}$ is an unbiased and consistent estimate of $\sigma^{2}$ Hogg and Craig (2019). Therefore, by replacing $\sigma$ by $s$, we can say that $Y=\frac{\langle X\rangle-m}{s/\sqrt{n}}$ is approximately standard normal for large $n$.
@@ V.1 Interval Estimation
$P(-y_{\epsilon}<Y<y_{\epsilon})$ $=P(-y_{\epsilon}<\frac{\langle X\rangle-m}{s/\sqrt{n}}<y_{\epsilon})$
@@
Now we can decide the photon count distribution as sub-Poissonian, or super-Poissonian, with confidence level $1-\epsilon$, if the estimated variance, $s\leq\langle x\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}$, or $s\geq\langle x\rangle+\frac{sy_{\epsilon}}{\sqrt{n}}$, i.e., $\frac{\langle x\rangle}{s}\geq 1+\frac{y_{\epsilon}}{\sqrt{n}}$, or $\frac{\langle x\rangle}{s}\leq 1-\frac{y_{\epsilon}}{\sqrt{n}}$ respectively, where $y_{\epsilon}$ is given by Eq. (13). That is, for some given $\epsilon$ and $y_{\epsilon}$ given by Eq. (13),
@@
with confidence level $1-\epsilon$.
@@
V.2 Hypothesis Testing
@@
$H_{0}:\ m=s^{2},\ H_{a}:\ m>s^{2}.$ (18)
For both tests, the test statistic would be $z=\frac{\langle x\rangle-s^{2}}{s/\sqrt{n}}$. For the first test,
@@
V.3 Goodness-of-fit Test
@@
For testing the quantumness in the generated random number, one may use the $\chi^{2}$ goodness-of-fit test Hogg and Craig (2019); Dodge (2008b) on the data generated by QRNGs as follows. Consider null and alternative hypotheses as
@@
is the well-known Gamma function Hogg and Craig (2019). The cumulative distribution function in this case is given by,
@@
As in most cases, the equality of the mean and the variance denotes Poisson distribution, for practical purposes, the last $\chi^{2}$ test may not be needed. We mention this test for the completeness of the discussion.
VI Detailed Photon Counting Statistics of Coherent Pulse-based QRNG
The statistics discussed in Eq. (11) of Section IV consider that all of the $n$ photons present within time $T$ would be detected in the same time-bin, which is a crude assumption for $n>1$. In this section, we perform a more fine-grained analysis of the first photon arrival time for coherent pulse-based QRNG devices. This analysis helps us to investigate the quality of QRNG in terms of several parameters, as elaborated in Section VII.
VI.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
Currently available photon arrival time-based QRNGs Qutools (2010); QNu Labs (2018) use coherent sources, which generate photons that follow Poissonian statistics. Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@ VI.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
However, we do not know the number of photons present within the period $T$. Therefore, the probability of detecting the first photon in the $i$-th bin is given by
@@
VI.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
Till now, we have considered all of the used optical devices as ideal. But in practice, any error in the source and in the attenuators affects photon power and changes the value of the mean photon number per unit time. Again, if the detector is not $100\%$ efficient, The first photon would be detected in the $i$-th time-bin, if at least one photon has been detected in $i$-th bin, and either (A) no photon has been reached at SPD before the $i$-th bin, or (B) $k$ number of photons have reached before the $i$-th bin, but none of them has been detected. Therefore, the probability of detecting the first photon in the $i$-th bin, when $n$ photons are present within $T$ time, is given by
$P_{d}(i;1st|n)$ $=\frac{\sum_{k=0}^{n-1}P(k,\frac{(i-1)T}{N})(1-d)^{k}\sum_{l=1}^{n-k}\{1-(1-d)^{l}\}P(l,\frac{T}{N})P(n-k-l,T-\frac{iT}{N})}{P(n,T)}$
$=\frac{\sum_{k=0}^{n-1}\frac{e^{-\mu T\frac{i-1}{N}}\left(\mu T\frac{i-1}{N}\right)^{k}}{k!}(1-d)^{k}\sum_{l=1}^{n-k}\{1-(1-d)^{l}\}\frac{e^{-\mu T/N}\left(\frac{\mu T}{N}\right)^{l}}{l!}\frac{e^{-\mu T\left(1-\frac{i}{N}\right)}(\mu T)^{n-k-l}\left(1-\frac{i}{N}\right)^{n-k-l}}{(n-k-l)!}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=n!\sum_{k=0}^{n-1}\frac{\left(\frac{(i-1)(1-d)}{N}\right)^{k}}{k!}\frac{1}{(n-k)!}\sum_{l=1}^{n-k}\binom{n-k}{l}\left[\left(\frac{1}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-k-l}-\left(\frac{1-d}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-k-l}\right]$
@@ VI.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
Figure 5 shows how $\mu Td$ affects the probability (29). Here we consider $N=256$. From this plot, it is clear that the randomness increases as $\mu Td$ decreases. Therefore, the main goal in designing a QRNG is
@@
However, this minimization is bounded by some constraints related to the generation rate, the cost and the timing error in registering the photon detection time. These constraints are discussed in Section VII.
@@
Since in the probability distribution of Eq. (VI.2), $\frac{T}{N}$ is the length of the time-bins, Eq. (32) is identical with Eq. (VI.2). Therefore, the probability mass function corresponding to the probability Eq. (32) is also given by Eq. (29) with $T=Nt_{i}$.
Therefore, these two types of QRNGs, perform similarly in practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (Ref. Eq. (11)). Since the distributions given by Eq. (29) are evaluated from the quantum properties of photons, this can be used to test the quantumness in a QRNG.
@@
$\max_{i,j\in S}|P(i)-P(j)|=\epsilon.$ (33)
Since the distribution (29) is a monotone decreasing function, it becomes $\epsilon$-random if and only if
$f(1)-f(N)=\epsilon$
$\implies\frac{e^{\frac{\mu Td}{N}}-1}{e^{\mu Td}-1}\left(e^{\left(1-\frac{1}{N}\right)\mu Td}-1\right)=\epsilon.$ (34)
Figure 6 is a visual representation of the Eq. (34). It clearly shows that the value of $\epsilon$ increases with $\mu Td$. That means, to generate numbers with $\epsilon$-randomness, we need to carefully choose the external reference cycle $T$, and configure the source, the attenuators and the detector so that relation (34) has been satisfied.
We have performed the NIST, Dieharder, AIS-31 and ENT tests on the data generated by simulating the probability distribution given by Eq. (29). These test results are provided in Section VIII.
VII Effect of $\mu,T,d$ on random number generation rate and cost
As discussed in Eq. (30) of Subsection VI.2, we have to minimize $\mu Td$ to achieve maximum randomness from the quantum source increases. In that discussion, we have considered the product $\mu Td$ as a single quantity. In this section, we separately discuss the effect of

### v4

@@ 3 QRNG based on Single-photon Source
Figure 1: A generic block diagram of the single-photon-based QRNG. Ideally, single-photon, coming from the source, are sent through a $/50$ 50 beam-splitter. Two SPDs ($D_{0},D_{1}$) are placed in two paths to detect the photon. Photon may take either path with probability $0.5$. The random bit would be $i$ if the detection happens at $D_{i}$. In practice, beam splitters and detectors cannot be error-free. However, for an $R/T$ beam-splitter, by choosing detectors as in Eq. (7), uniform random numbers can be generated.
3.1 Effect of Imperfect Devices
If the photon source for this QRNG is not a perfect single-photon source, it may happen that at the same time some photons are transmitted through the beam-splitter and some are reflected. Then, both the detectors may detect photons simultaneously, and the random number cannot be generated. To ensure whether the source is a single-photon source or not, one may perform the first step of the two-fold statistical method mentioned in Section 5. If this method reveals that the distribution is not sub-Poissonian, then the source cannot be a single-photon source, and that source cannot be used in this QRNG.
Each bit of the random number generated with the above single-photon-based QRNG has probability $\frac{1}{2}$ of being 00 or $1$ when all the underlying hardware components such as source, beam-splitter, SPD, are ideal. So, ideally, it produces uniform random numbers. However, any defect in the beam-splitter or the detectors may change this probability. Let us consider an $R/T$ beam-splitter and the detectors $D_{i}$ with detection efficiency $d_{i}\in(0,1]$. Suppose, $D_{0}$ is placed on the path of the transmitted light, and $D_{1}$ is placed on the path of the reflected light. Then the probability of 00 in the generated random number, that is, the probability of detecting a photon by detector $D_{0}$ would be
@@ 3.1 Effect of Imperfect Devices
Therefore, to generate random numbers, we require,
$P(D_{0})=P(D_{1})\implies d_{0}T=d_{1}R.$ (7)
If $d_{0}T\neq d_{1}R$, the generated random number would not be uniformly random. To test this randomness, statistical tests like NIST, Dieharder, AIS-$31$, ENT etc. may be used. See Section 8 for a discussion on these.
4 Existing QRNG Models based on Coherent Pulse
It is well known that producing single photon states is experimentally challenging [54, 55]. However, different experimental approaches [56, 57, 58, 59, 60, 61, 62, 63] have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval [96]. In this process, the photons follow Poissonian statistics. A generic block diagram of the photon arrival time-based QRNG using photons from coherent sources is given in Fig. 2.
@@ 4 Existing QRNG Models based on Coherent Pulse
Figure 2: A basic block diagram of the photon arrival time-based QRNG using weak coherent pulses. Coherent pulses, coming from the source, are sent through attenuators to reduce their power so that the probability of having multiple photons within a predefined time interval becomes negligible [96]. A SPD detects these photons and the detection time (photon arrival time) has been registered. The Field Programmable Gate Array (FPGA) provides random numbers based on the registered time stamps.
@@
There are different methods to produce random numbers from the registered times. Time diagrams for these methods are shown in Fig. 3. In Fig. 3(I), if the registered times of photon arrivals are $\{t_{0},t_{1},\dots,t_{2i-2},t_{2i-1},t_{2i},\dots\}$, the non-overlapping consecutive time intervals, i.e., $\Delta t_{2i-1}=t_{2i-1}-t_{2i-2},\Delta t_{2i}=t_{2i}-t_{2i-1}$, are used to generate the $i$-th bit of the random numbers as follows.
[…]
But in this process [13], the random number generation rate is very slow. The length of the random bit-string is half of the photon count. Comparatively faster processes have been developed in the literature [12, 15, 14, 16, 17, 17]. In Fig. 3(II), the time is divided into time-bins, each having equal length. If a photon is detected in the $i$-th time-bin, $i-1$ is taken as the random number. A new cycle of the bins starts after the detector has a fixed sleep time. Since the photon counting process is a Poisson process, the probability, that the waiting time ($t_{w}$) between the arrivals of two consecutive photons is greater than $t$, is given by
@@
Therefore, the random number generated with this procedure follows exponential distribution [15, 14, 16, 17]. Another QRNG scheme is proposed in [12], as depicted in Fig. 3(III), using external time reference. This external time is divided into cycles of length $T$, which is smaller than the dead time of the SPD. The time $T$ is further divided into $N$ equal subintervals, called the time-bins $\tau_{1},\tau_{2},\dots,\tau_{N}$. In each cycle, if the first photon is detected in the $i$-th time-bin, $i-1$ would be the corresponding random number. SPD would be in sleep mode for the remaining time of the cycle.
@@
5 Our Proposal of Two-fold Statistical Method for Testing Quantumness
In Section 2 we have discussed that for the Poissonian distribution, the mean of the photon-counting distribution is the same as its variance. However, it is interesting to note that mean $=$ variance does not always imply that the distribution is Poisson. For example, consider the probability distribution
@@ 5 Our Proposal of Two-fold Statistical Method for Testing Quantumness
This motivates us to propose a two-fold method to test the quantumness of the photon source as an answer to question Q1. First, a statistical test is performed to verify whether the mean is equal to or less than or greater than the variance. Note that any statistical validation of equality is not exact, but rather approximate in practice. Thus, if the above test validates the mean and the variance to be equal, we propose to perform a second test to validate that the distribution is indeed Poisson. The statistical tools have been used in this section is discussed in Appendix C.
5.1 First Phase: Interval Estimation
Since Poissonian and super-Poissonian statistics are explainable using classical theory, only sub-Poissonian distribution can guarantee the quantumness of a quantum device. We suggest a procedure to decide whether the underlying distribution is sub-Poissonian or not by estimating the confidence interval for the mean of the underlying distribution as follows.
Suppose, the number of photon counts within a fixed time is given as a sample of size $n$. Then by the central limit theorem, $\frac{\langle X\rangle-m}{\sigma/\sqrt{n}}$ is approximately standard normal for large $n$  [97], where $\langle X\rangle$ is the random variable corresponding to the sample mean $\langle x\rangle$, and $m,\sigma^{2}$ are mean and variance of the actual photon-count distribution. Let the sample variance be given by $S^{2}$. Then
$s^{2}=\frac{n}{n-1}S^{2}$ (14)
is an unbiased and consistent estimate of $\sigma^{2}$ [97]. Therefore, by replacing $\sigma$ by $s$, we can say that $Y=\frac{\langle X\rangle-m}{s/\sqrt{n}}$ is approximately standard normal for large $n$.
@@ 5.1 First Phase: Interval Estimation
$P(-y_{\epsilon}<Y<y_{\epsilon})$
$=P(-y_{\epsilon}<\frac{\langle X\rangle-m}{s/\sqrt{n}}<y_{\epsilon})$
@@
Now we can decide the photon count distribution as sub-Poissonian, or super-Poissonian, with confidence level $1-\epsilon$, if the estimated variance, $s^{2}\leq\langle x\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}$, or $s^{2}\geq\langle x\rangle+\frac{sy_{\epsilon}}{\sqrt{n}}$, i.e., $\frac{\langle x\rangle}{s^{2}}\geq 1+\frac{y_{\epsilon}}{s\sqrt{n}}$, or $\frac{\langle x\rangle}{s^{2}}\leq 1-\frac{y_{\epsilon}}{s\sqrt{n}}$ respectively, where $y_{\epsilon}$ is given by Eq. (15). That is, for some given $\epsilon$ and $y_{\epsilon}$ given by Eq. (15),
@@
distribution with confidence level $1-\epsilon$.
@@
5.2 Hypothesis Testing as an Alternative to Interval Estimation
@@
$H_{0}:\ m=s^{2},\ H_{a}:\ m>s^{2},$ (20)
where $m$ is the mean of the actual distribution and $s^{2}$ is given by Eq. (14). For both tests, the test statistic would be $z=\frac{\langle x\rangle-s^{2}}{s/\sqrt{n}}$. For the first test,
@@
5.3 Second Phase: Goodness-of-fit Test
@@
For testing the quantumness in the generated random number, one may use the $\chi^{2}$ goodness-of-fit test [97, 98] on the data generated by QRNGs as follows. Consider null and alternative hypotheses as
@@
is the well-known Gamma function [97]. The cumulative distribution function in this case is given by,
@@
As in most cases, the equality of the mean and the variance denotes Poisson distribution, for practical purposes, the last $\chi^{2}$ test may not be needed. We mention this test to ensure the completeness of the discussion.
Note that if synthetic data is generated via simulation according to the actual photon statistics distribution, our two-fold method might inaccurately classify this data as coming from a sub-Poissonian quantum source. We do not assert that our two-fold method can classify any off-the-shelf black-box random number generator. Rather, we claim that, given the internal parameters and intermediate distributions from a real device, our two-fold method provides a more rigorous validation of whether the optical source operates in the classical or quantum regime compared to merely comparing the exact equality of mean and variance, thus confirming the quantum nature of the generated random numbers.
6 Detailed Photon Counting Statistics of Coherent Pulse-based QRNG
The statistics discussed in Eq. (12) of Section 4 consider that all of the $n$ photons present within time $T$ would be detected in the same time-bin, which is a crude assumption for $n>1$. In this section, we perform a more fine-grained analysis of the first photon arrival time for coherent pulse-based QRNG devices. This analysis helps us to investigate the quality of QRNG in terms of several parameters, as elaborated in Section 7.
6.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
Currently available photon arrival time-based QRNGs [37, 38] use coherent sources, which generate photons that follow Poissonian statistics. Here, a random number is generated based on the detection time of the first photon in each time cycle $T$. The probability of detecting the first photon in $i$-th bin, when $n$ photons are present within $T$ time, is given by
@@ 6.1 Time-bin Statistics for Poissonian Photon with Perfect Devices
This was also observed by Nie et al. [17, Eq. (2)]. However, we do not know the number of photons present within the period $T$. Therefore, the probability of detecting the first photon in the $i$-th bin is given by
@@
6.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
Until now, we have considered all of the used optical devices as ideal. But in practice, any error in the source and in the attenuators affects photon power and changes the value of the mean photon number per unit time. Again, if the detector is not $100\%$ efficient, the first photon would be detected in the $i$-th time-bin, if at least one photon has been detected in $i$-th bin, and either (A) no photon has been reached at SPD before the $i$-th bin, or (B) $k$ number of photons have reached before the $i$-th bin, but none of them has been detected. Therefore, the probability of detecting the first photon in the $i$-th bin, when $n$ photons are present within $T$ time, is given by
$P_{d}(i;1st|n)$ $=\frac{\sum\limits_{k=0}^{n-1}P(k,\frac{(i-1)T}{N})(1-d)^{k}\sum\limits_{l=1}^{n-k}\{1-(1-d)^{l}\}P(l,\frac{T}{N})P(n-k-l,T-\frac{iT}{N})}{P(n,T)}$
$=\frac{\sum\limits_{k=0}^{n-1}\frac{e^{-\mu T\frac{i-1}{N}}\left(\mu T\frac{i-1}{N}\right)^{k}}{k!}(1-d)^{k}\sum\limits_{l=1}^{n-k}\{1-(1-d)^{l}\}\frac{e^{-\mu T/N}\left(\frac{\mu T}{N}\right)^{l}}{l!}\frac{e^{-\mu T\left(1-\frac{i}{N}\right)}(\mu T)^{n-k-l}\left(1-\frac{i}{N}\right)^{n-k-l}}{(n-k-l)!}}{\frac{e^{-\mu T}(\mu T)^{n}}{n!}}$
$=n!\sum_{k=0}^{n-1}\frac{\left(\frac{(i-1)(1-d)}{N}\right)^{k}}{k!(n-k)!}\sum_{l=1}^{n-k}\binom{n-k}{l}\left[\left(\frac{1}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-k-l}-\left(\frac{1-d}{N}\right)^{l}\left(1-\frac{i}{N}\right)^{n-k-l}\right]$
@@ 6.2 Time-bin Statistics for Poissonian Photon with Inperfect Devices
Note that $f(1)$ is the probability $P_{1}$ in Ref. [17, Eq. (4)]. As their aim was to find the minimum entropy for deciding the amount of post-processing, they only derived an upper bound to $P_{1}$. However, we aim to reduce the amount of post-processing. Therefore, we have derived the complete probability mass function to see how the device parameters, $\mu,T$ and $d$ affect the distribution. We also discuss the effect of all the parameters separately in Section 7. Figure 5 shows how $\mu Td$ affects the probability (31). Here we consider $N=256$. From this plot, it is clear that the uniformity increases as $\mu Td$ decreases.
Although this uniformity can be increased by post-processing the raw data, as mentioned above, this reduces the random number generation rate and increases the possibility of side-channel attack. Since here we want to minimize the post-processing, the main goal in designing a QRNG is
@@
However, this minimization is bounded by some constraints related to the generation rate, the cost and the timing error in registering the photon detection time. These constraints are discussed in Section 7.
@@
Since in the probability distribution of Eq. (6.2), $\frac{T}{N}$ is the length of the time-bins, Eq. (34) is identical with Eq. (6.2). Therefore, the probability mass function corresponding to the probability Eq. (34) is also given by Eq. (31) with $T=Nt_{i}$.
Therefore, these two types of QRNGs, perform similarly in practical scenarios. However, if we can ensure that the source is an actual single-photon source, then the perfect random number can be generated using an external time reference with probability distribution $P_{d}(i)=\frac{d}{N}$ (Ref. Eq. (12)). Since the distributions given by Eq. (31) are evaluated from the quantum properties of photons, this can be used in the goodness-of-fit test mentioned in Subsection 5.3 to test the quantumness in the raw data generated by a QRNG.
@@
$\max_{V\subseteq S}\left|\sum_{i\in V}\left(P(i)-\frac{1}{|S|}\right)\right|=\frac{1}{2}\sum_{i\in S}\left|P(i)-\frac{1}{|S|}\right|$ $=\epsilon.$ (35)
Therefore, the distribution (31) becomes $\epsilon$-random if and only if
$\frac{1}{2}\sum_{i\in S}\left|f(i)-\frac{1}{N}\right|=\epsilon.$ (36)
Figure 6 is a visual representation of the Eq. (36) with $N=256$. It clearly shows that the value of $\epsilon$ increases with $\mu Td$. That means, to avoid post-processing while generating random numbers with $\epsilon$-randomness for a given $\epsilon$, one needs to carefully choose the external reference cycle $T$, and configure the source, the attenuators and the detector so that relation (36) has been satisfied.
We have performed the NIST, Dieharder, AIS-31 and ENT tests on the data generated by simulating the probability distribution given by Eq. (31). These test results are provided in Section 8.
7 Effect of $\mu,T,d$ on random number generation rate and cost
As discussed in Eq. (32) of Subsection 6.2, we have to minimize $\mu Td$ to achieve maximum randomness from the quantum source increases. In that discussion, we have considered the product $\mu Td$ as a single quantity. In this section, we separately discuss the effect of